OpenStudy (abbie):

solve for x. -3x+7y=2 -2x-3y=-4 A. x=-22/23 B. x=22/23 C. x=2/23 D. x=-2/23

OpenStudy (abbie):

@agent0smith can you help me?

OpenStudy (abbie):

@skullpatrol

OpenStudy (abbie):

@KamiBug help?

OpenStudy (kamibug):

I know I did this in 8th grade but I don't remember it, sorry!

OpenStudy (texaschic101):

does it have to be done a certain way...elimination, substitution, matrix...or does it matter ?

OpenStudy (texaschic101):

this is elimination -3x + 7y = 2 --->(3) -2x - 3y = -4 --->(7) -------------------- -9x + 21y = 6 (result of multiplying by 3) -14x - 21y = -28 (result of multiplying by 7) -------------------add -23x = -22 x = 22/23